OpenStudy (romantic_ch3micals):

Help? Medal. and...*sighs* I'll fan you The area of a the rectangular painting is given by the trinomial x^2+2x-15. What are possible dimensions? Use factoring

OpenStudy (romantic_ch3micals):

...erm anyone?

OpenStudy (mathmate):

See: http://openstudy.com/study#/updates/5563dadee4b07daafb7ea498 At the end

OpenStudy (romantic_ch3micals):

...erm anyone?

OpenStudy (romantic_ch3micals):

...erm. I need help on this on. Besides, I don't think the teacher taught us this way.

OpenStudy (mathmate):

Well, you'd look for two factors of -15 that adds up to +2 Either you guess randomly, or you make a table: -15*1 = -15 -15+1 = -14 -5*3=-15 -5+3 = -2 but 5-3=+2, so try 5*(-3)=-15 5-3 = +2 So x^2+2x-15 = (x+5)(x-3) Or you can use the method in the link if the coefficient is not one.